BWP projects sales of 100000 units next year at an average price of $50 per unit. Variable costs are estimated at 40% of revenue, and fixed costs will be $2.4 million. BWP has $1 million in bonds outstanding, on which it pays 7.5%, and its marginal tax rate is 40%. There are 100000 shares of stock outstanding which trade at their book value of $30. Compute BWP's contribution, contribution margin, net income, DOL, and EPS. Round the answers to two decimal places. Enter your net income answer in whole dollars. For example, an answer of $1.2 million should be entered as 1,200,000, not 1.2.

Contribution per unit $
Contribution Margin      %
Net Income $
DOL     
EPS
Answer #1
Average sale price per unit $50
Variable costs @ 40% (50*40%) $20
Contribution margin per unit (50-20) $30
Contribution margin ratio (30/50) 60%
Sales revenue 5000000
Variable costs (5,000,000*40%) 2000000
Contribution margin 3000000
Fixed costs 2400000
EBIT 600000
Interest (1,000,000*7.5%) 75000
EBT 525000
Taxes @ 40% 210000
Net Income 315000
DOL = Contribution Margin/EBIT
DOL = 3,000,000/600,000 = 5
EPS = Net income/ Shares outstanding
EPS = 315,000/100,000 = 3.15 per share